OpenStudy (anonymous):

A spherical balloon bursts when the surface area(S) exceeds 500 cm2. If the radius changes at a rate of 23 cm/s, what is the rate of change of volume at the instant the balloon bursts? ... I have no idea how to figure this out, please help guys :/

OpenStudy (cwrw238):

dV/ dt = dV/dr * dr / dt dr/dt = 23 can you find an expression for dV/dr?

OpenStudy (anonymous):

I can find dV, how would i find dr though?

OpenStudy (cwrw238):

volume of a sphere V = (4/3) pi r^3 now find the derivative with respect to r i'm assuming you know some calculus

OpenStudy (anonymous):

4(pi)r^2, right?

OpenStudy (cwrw238):

yes - now what is 4 pi r^2 ?

OpenStudy (cwrw238):

- its the formula for what/

OpenStudy (anonymous):

surface area

OpenStudy (cwrw238):

ok - so now its easy you know that dV/dr = 4 pi r^2 = 500 and dr/dt = 23

OpenStudy (cwrw238):

so multiply

OpenStudy (anonymous):

so 11500cm^3/s is the final answer?

OpenStudy (anonymous):

thanks :D that was easier than i thought

OpenStudy (cwrw238):

yw